OpenStudy (anonymous):

f(x)=x^(2)*(sin(1/x)) is it differentiable at x=0

OpenStudy (anonymous):

a function being differentiable at a given point means the derivative is continuous at that point. the derivative of f is 2xsin(1/x)-cos(1/x). it's fairly straightforward to check continuity at 0.

OpenStudy (anonymous):

thanks cnknd!! but i have doubt that does cos(1/x) also wobbles between 0 ,-1 and 1 like sin1/x

OpenStudy (anonymous):

ok, lets break that thing into 2 parts first: 2xsin(1/x), is this continuous at 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2nd part, cos(1/x): is this continuous at 0?

OpenStudy (anonymous):

remember what it means to be continuous: the left and right limits exist and are equal

OpenStudy (anonymous):

cos1/x is similar to sin1/x as x approaches to 0 i.e it keeps oscillating between 0 ,1 ,-1 as x approaches 0

OpenStudy (anonymous):

well the answer here is that neither the left nor right limits exist for cos(1/x) as x goes to 0 (showing why this is true is a bit more involved, let me know if u want to go through the rigor)

OpenStudy (anonymous):

yes please that will be helpful

OpenStudy (anonymous):

ok, so let's look at the right limit of cos(1/x) as x -> 0 i'll first make a change of variables: let y = 1/x

OpenStudy (anonymous):

so this limit now becomes: limit of cos(y) as y -> positive infinity

OpenStudy (anonymous):

well, i havent specified how y is going to infinity. perhaps it goes as the following sequence: 1, 2, 3, 4, ....; or this sequence: 3, 5, 7, 9, 11, ...

OpenStudy (anonymous):

if i can identify 2 ways (2 different sequences) for y to approach infinity, but the 2 paths would give different limits of the function cos(y), then it doesn't make sense to say that the limit of the function exists

OpenStudy (anonymous):

following me so far?

OpenStudy (anonymous):

i dint get that 2 sequences

OpenStudy (anonymous):

a sequence here just means how y is going to infinity.

OpenStudy (anonymous):

so here's one way for y to approach infinity: y = 0, 2pi, 4pi, 6pi, ..... and so the value of cos(y), as y approaches infinity in this path, goes as: cos(y) = 1, 1, 1, 1, ... so without going further, have i convinced you that the limit of cos(y) is 1?

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

well here's another sequence for y: y = pi, 3pi, 5pi, .... so we have: cos(y) = -1, -1, -1, ... and it looks like the limit here (for the same function) is -1. the limit can't possibly be -1 and 1 at the same time. therefore it's undefined

OpenStudy (anonymous):

y dint u take npi/2 for various values of n in consideration?

OpenStudy (anonymous):

i could, and that would give me a limit of 0

OpenStudy (anonymous):

still the same conclusion though, the limit can't be -1, 0, and 1 at the same time, so it does not exist

OpenStudy (anonymous):

ok thanks cnknd!!! that means f(x) is not differentiable at x= 0

OpenStudy (anonymous):

that is correct.

OpenStudy (anonymous):

in one of the class notes i saw that lhl, f'(0) =limf(0-h)-f(0)/h=h*sin(1/h) =0 ;h --> 0

OpenStudy (anonymous):

and rhl, f'(0)=lim f(0+h) - f(0)/h =h* sin1/h = 0

OpenStudy (anonymous):

lhl= rhl thereby showing that f'(x)is continuous at x= 0 , i cant find the mistake

OpenStudy (anonymous):

i have no idea how they did this step: (f(0+h)-f(0))/h = h*sin(1/h)

OpenStudy (anonymous):

they have shown f(0+h)= h^2 * sin(1/h) aand f(0)= 0 and then divided by h

OpenStudy (anonymous):

o.o sorry i think i was wrong... the wikipedia page has this function as an example of differentiable but not smooth: http://en.wikipedia.org/wiki/Differentiable_function i apologize for misleading you

OpenStudy (anonymous):

but even in your approach i didnt find anything wrong!! i am really confused

OpenStudy (anonymous):

oh i went wrong in the definition: i thought it was asking for smooth (i.e. differentiable, and derivative is continuous). where as it was simply asking for differentiability

OpenStudy (anonymous):

being differentiable and derivative is continuous? are they both different?

OpenStudy (anonymous):

the distinction here is not a place in mathematics that I've looked at in detail. but wikipedia says a function can have certain kinds of discontinuity in the derivative, and still be differentiable: http://en.wikipedia.org/wiki/Differentiable_function#Differentiability_classes

OpenStudy (anonymous):

so to test differentiability what should be my approach for any given f(x)

OpenStudy (anonymous):

ur lecture notes would be correct for simple differentiability i.e. use the definition of the derivative.

OpenStudy (anonymous):

i still dont understand that if we take derivative of f(x) and try to check its continuity it is discontinuous....and yet it is continuous using the defintion of derivative why two different answer?

OpenStudy (anonymous):

oh the definition of the derivative doesnt say that the derivative is continuous

OpenStudy (anonymous):

the limits in the definition of the derivative is telling you whether or not your approximations of the derivative is continuous (as they get more and more accurate to the true derivative)

OpenStudy (anonymous):

but y two different answer...because even the former approach seems to be correct

OpenStudy (anonymous):

well it's not asking for whether or not the derivative is continuous, it's just asking for whether or not the derivative exists

OpenStudy (anonymous):

ok to test differentiability i should use limf(x+h)-f(h)/h as h --> 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks cnknd!!